Solve the equation 3log5 x - log5 4 = log5 16
Answer by consc198(59)

3log5 x - log5 x 4 = log5 x 16
2.096910013 x -0.698970004 x 4 = 0.698970004 x 16
2.096910013 x -2.795880016 = 11.18352006
-5.862708801 = 11.18352006
11.18352006 + 5.862708801
Answer = 17.04622886